# Hybrid Bernstein Normalizing Flows for Flexible Multivariate Density Regression with Interpretable Marginals

### v1

@@ Results
To assess whether the model captures the marginal distribution, Hothorn et al. [2014] suggest using quantile-quantile (Q-Q) plots to verify $\mathbf{W}$ follows the base distribution $F_{W}=\mathcal{N}(0,1)$. Figure 6 shows the empirical quantile function of marginally normalized samples $\mathbf{W}$ from the validation set plotted against the quantile function of a standard normal distribution evaluated at 200 equidistant probabilities. The results suggest a good fit of the transformation $H_{1}$, with slight underestimations of the marginals.
@@ A.1 Bernstein Polynomials
where $\operatorname{Be}_{i}(z)$ is the density of a Beta distribution with parameters $i+1$ and $M-i+1$, and $\vartheta_{0},\ldots,\vartheta_{M}$ are the Bernstein coefficients Farouki [2012]. As $M$ increases, Bernstein polynomials become a universal approximator for smooth functions in $[0,1]$ Farouki and Rajan [1988]. In practice, $M\geq 10$ is often sufficient Hothorn et al. [2018].
Bernstein polynomials are defined for values of $z$ within the range $[0,1]$. Outside this interval, linear extrapolation is performed. To guarantee invertibility, transformation functions must be bijective, achieved by strict monotonicity. When using Bernstein polynomials, monotonicity is enforced by constraining the Bernstein coefficients $\vartheta_{0},\ldots,\vartheta_{M}$ to be increasing. This is achieved by recursively applying a strictly positive function like $\operatorname{softplus}$ to an unconstrained vector $\tilde{\vartheta}_{0},\ldots,\tilde{\vartheta}_{M}$, such that $\vartheta_{0}=\tilde{\vartheta}_{0}$ and $\vartheta_{k}=\vartheta_{k-1}+\operatorname{softplus}(\tilde{\vartheta}_{k})$ for $k=1,\ldots,M$. Parameter estimation can depend on initialization, as $\vartheta_{0}$ is directly derived from unconstrained parameters. Monotonicity of $h$ is enforced by constraining the coefficients to be increasing, e.g. by $\vartheta_{0}=\tilde{\vartheta}_{0}$ and $\vartheta_{k}=\vartheta_{k-1}+\operatorname{softplus}(\tilde{\vartheta}_{k})$ for $k=1,\ldots,M$ Sick et al. [2021]. The inverse transformation can be found using a root-finding algorithm Chandrupatla [1997].
@@ A.2 Maximum Likelihood Estimation for Parameter Inference
Parameter estimation usually involves minimizing a divergence between the true distribution $p_{y}(y)$ and the transformation model $p_{y}(\mathbf{y}|\theta)$ regarding the model’s parameters $\theta$ [Papamakarios et al., 2021]. Since $p_{y}(y)$ is unknown, we minimize the negative log-likelihood of the empirical distribution $p_{\mathcal{D}}$ obtained from a finite set $\mathcal{D}$ of $N$ i.i.d. observations $\mathbb{Y}=\{\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\}$:
@@
which is equivalent to minimizing the KL divergence between $p_{\mathcal{D}}$ and the flow-based model $p_{y}(y;\theta)$ [Papamakarios et al., 2021].
@@ A.3 Multivariate Conditional Transformation Models
MCTM [Klein et al., 2019] use element-wise transformations $\tilde{h}_{j}\left(y_{j}\right),j=1,\ldots,J$ and a linear triangular $(J\times J)$ matrix $\Lambda$ to model correlations:
@@ A.4 Autoregressive Transformation Models
This is implemented as $z_{i}=h_{i}(y_{i},\theta_{i})$ with $\theta_{i}=c(\mathbf{y}_{<i})$, where the conditioner $c_{i}$, often a neural network, captures dependencies and enforces the autoregressive property [Papamakarios et al., 2021]. New samples are obtained via the inverse transformation $y_{i}=T_{i}^{-1}(z_{i},\theta_{i})$.
Two common architectures for the conditioner are coupling layers and masked autoregressive networks. Coupling flows [Dinh et al., 2017] split the response $\mathbf{y}\in\mathbb{R}^{D}$ into subsets $(\mathbf{y}_{A},\mathbf{y}_{B})\in(\mathbb{R}^{d},\mathbb{R}^{D-d})$, with one subset conditioning the transformation of the other:
@@
MADE [Germain et al., 2015], as used in MAFs [Papamakarios et al., 2017] and inverse autoregressive flows (IAFs) [Kingma and Ba, 2017], generalizes this using masked neural networks to enforce autoregressive constraints.
In MADE, the output $d$ depends only on inputs $\mathbf{y}_{<d}$. This is achieved by element-wise multiplying weight matrices by binary masks, zeroing connections that violate the autoregressive property. Whether the masked network uses $\mathbf{y}$ or the latent representation $\mathbf{z}$ as input affects only whether inference or sampling is iterative [Papamakarios et al., 2021].
@@ A.5 Structured Additive Predictors for Enhanced Flexibility
Structured Additive Predictors [SAP; Fahrmeir et al., 2013, 2009] allow for (non-)linear effects, interactions, and other structured terms. An exemplary predictor is given by
@@ B.3 Interpretability of Covariate Effects
Interpretable covariate effects are important in many regression tasks. CTMs offer a balance between interpretability and flexibility [Hothorn et al., 2014]. The base distribution and transformation function parameterization determine the interpretability.
@@ B.4 Interpretable Covariate Effect on the Dependency Structure in MCTMs
As described in Klein et al. [2019], the covariate effect on the dependency in the MCTM model can be interpreted as Spearman’s rank correlation from the covariance matrix
@@ B.5 Benchmark Datasets
We evaluate our method on five common benchmark datasets: POWER [^5] https://archive.ics.uci.edu/dataset/235, GAS [^6] https://archive.ics.uci.edu/dataset/322. HEPMASS [^7] https://archive.ics.uci.edu/dataset/347, MINIBOONE [^8] https://archive.ics.uci.edu/dataset/199/, BSDS300999https://www2.eecs.berkeley.edu/Research/Projects/CS/vision/bsds/. We follow the preprocessing steps from Papamakarios et al. [2017] [^10] https://github.com/francois-rozet/uci-datasets.

### v2

@@ Results
To assess whether the model captures the marginal distribution, Hothorn2014 suggest using quantile-quantile (Q-Q) plots to verify $\mathbf{W}$ follows the base distribution $F_{W}=\mathcal{N}(0,1)$. Figure 6 shows the empirical quantile function of marginally normalized samples $\mathbf{W}$ from the validation set plotted against the quantile function of a standard normal distribution evaluated at 200 equidistant probabilities. The results suggest a good fit of the transformation $H_{1}$, with slight underestimations of the marginals.
@@ Acknowledgements.
[…]
Hybrid Bernstein Normalizing Flows for Flexible Multivariate Density Regression with Interpretable Marginal
(Supplementary Material)
@@ A.1 Bernstein Polynomials
where $\operatorname{Be}_{i}(z)$ is the density of a Beta distribution with parameters $i+1$ and $M-i+1$, and $\vartheta_{0},\ldots,\vartheta_{M}$ are the Bernstein coefficients Farouki2012. As $M$ increases, Bernstein polynomials become a universal approximator for smooth functions in $[0,1]$ Farouki1988. In practice, $M\geq 10$ is often sufficient Hothorn2018.
Bernstein polynomials are defined for values of $z$ within the range $[0,1]$. Outside this interval, linear extrapolation is performed. To guarantee invertibility, transformation functions must be bijective, achieved by strict monotonicity. When using Bernstein polynomials, monotonicity is enforced by constraining the Bernstein coefficients $\vartheta_{0},\ldots,\vartheta_{M}$ to be increasing. This is achieved by recursively applying a strictly positive function like $\operatorname{softplus}$ to an unconstrained vector $\tilde{\vartheta}_{0},\ldots,\tilde{\vartheta}_{M}$, such that $\vartheta_{0}=\tilde{\vartheta}_{0}$ and $\vartheta_{k}=\vartheta_{k-1}+\operatorname{softplus}(\tilde{\vartheta}_{k})$ for $k=1,\ldots,M$. Parameter estimation can depend on initialization, as $\vartheta_{0}$ is directly derived from unconstrained parameters. Monotonicity of $h$ is enforced by constraining the coefficients to be increasing, e.g. by $\vartheta_{0}=\tilde{\vartheta}_{0}$ and $\vartheta_{k}=\vartheta_{k-1}+\operatorname{softplus}(\tilde{\vartheta}_{k})$ for $k=1,\ldots,M$ Sick2021. The inverse transformation can be found using a root-finding algorithm Chandrupatla1997.
@@ A.2 Maximum Likelihood Estimation for Parameter Inference
Parameter estimation usually involves minimizing a divergence between the true distribution $p_{y}(y)$ and the transformation model $p_{y}(\mathbf{y}|\theta)$ regarding the model’s parameters $\theta$ (Papamakarios2021). Since $p_{y}(y)$ is unknown, we minimize the negative log-likelihood of the empirical distribution $p_{\mathcal{D}}$ obtained from a finite set $\mathcal{D}$ of $N$ i.i.d. observations $\mathbb{Y}=\{\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\}$:
@@
which is equivalent to minimizing the KL divergence between $p_{\mathcal{D}}$ and the flow-based model $p_{y}(y;\theta)$ (Papamakarios2021).
@@ A.3 Multivariate Conditional Transformation Models
MCTM (Klein2022) use element-wise transformations $\tilde{h}_{j}\left(y_{j}\right),j=1,\ldots,J$ and a linear triangular $(J\times J)$ matrix $\Lambda$ to model correlations:
@@ A.4 Autoregressive Transformation Models
This is implemented as $z_{i}=h_{i}(y_{i},\theta_{i})$ with $\theta_{i}=c(\mathbf{y}_{<i})$, where the conditioner $c_{i}$, often a neural network, captures dependencies and enforces the autoregressive property (Papamakarios2021). New samples are obtained via the inverse transformation $y_{i}=T_{i}^{-1}(z_{i},\theta_{i})$.
Two common architectures for the conditioner are coupling layers and masked autoregressive networks. Coupling flows (Dinh2017) split the response $\mathbf{y}\in\mathbb{R}^{D}$ into subsets $(\mathbf{y}_{A},\mathbf{y}_{B})\in(\mathbb{R}^{d},\mathbb{R}^{D-d})$, with one subset conditioning the transformation of the other:
@@
MADE (Germain2015), as used in MAFs (Papamakarios2018) and inverse autoregressive flows (IAFs) (Kingma2016), generalizes this using masked neural networks to enforce autoregressive constraints.
In MADE, the output $d$ depends only on inputs $\mathbf{y}_{<d}$. This is achieved by element-wise multiplying weight matrices by binary masks, zeroing connections that violate the autoregressive property. Whether the masked network uses $\mathbf{y}$ or the latent representation $\mathbf{z}$ as input affects only whether inference or sampling is iterative (Papamakarios2021).
@@ A.5 Structured Additive Predictors for Enhanced Flexibility
Structured Additive Predictors (SAP; Fahrmeir2013; Fahrmeir2009) allow for (non-)linear effects, interactions, and other structured terms. An exemplary predictor is given by
@@ B.3 Interpretability of Covariate Effects
Interpretable covariate effects are important in many regression tasks. CTMs offer a balance between interpretability and flexibility (Hothorn2014). The base distribution and transformation function parameterization determine the interpretability.
@@ B.4 Interpretable Covariate Effect on the Dependency Structure in MCTMs
As described in Klein2022, the covariate effect on the dependency in the MCTM model can be interpreted as Spearman’s rank correlation from the covariance matrix
@@ B.5 Benchmark Datasets
We evaluate our method on five common benchmark datasets: POWER [^5] https://archive.ics.uci.edu/dataset/235, GAS [^6] https://archive.ics.uci.edu/dataset/322. HEPMASS [^7] https://archive.ics.uci.edu/dataset/347, MINIBOONE [^8] https://archive.ics.uci.edu/dataset/199/, BSDS300999https://www2.eecs.berkeley.edu/Research/Projects/CS/vision/bsds/. We follow the preprocessing steps from Papamakarios2018101010https://github.com/francois-rozet/uci-datasets.